OpenStudy (anonymous):

Help pretty please !!! the selling price of a stereo is $30 less than 3 times the cost. What percent of the selling price is that profit?

OpenStudy (anonymous):

let cost be c

OpenStudy (anonymous):

okay do 3*c-30

OpenStudy (anonymous):

so

OpenStudy (anonymous):

so profit would be (3c-30)-c

OpenStudy (anonymous):

Profit /selling price * 100

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

so if the selling price is 390 it would be

OpenStudy (anonymous):



OpenStudy (anonymous):

3*c-30=390 find c from this

OpenStudy (anonymous):

140

OpenStudy (primeralph):

pretty please? lol...well you're on track

OpenStudy (anonymous):

c=120

OpenStudy (anonymous):

if you plug 120 in for c it would be 330 not 390...

OpenStudy (anonymous):

i though c= 140

OpenStudy (anonymous):

(140*3)-30=390 420-30=390 390=390